Julie is 4 times older than Aaron. In 10 years, Julie will be twice as old as Aaron. How old are Julie and Aaron?
Thank you!:o)

Answer by checkley77(12844)

J=A+4
J+10=2(A+4)
A+4+10=2A+8
A+14=2A+8
A-2A=8-14
-A=-6
A=6 YEARS OLD FOR AARON.
J=6+4=10 YEARS OLD FOR JULIE.
PROOF:
10+10=2(6+4)
20=2*10
20=20